In 2004, the Diabetic Express charged $27.06 for a vial of Humulin insulin and $34.39 for a vial of Novolin Velosulin insulin. If a total of $1565.57 was collected for 50 vials of insulin, how many vials of each type were sold?
Answer by checkley77(12844)

34.39x+27.06(50-x)=1,565.57
34.39x+1,353-27.06x=1,565.57
7.33x=1,565.57-1,353
7.33x=212.57
x=212.57/7.33
x=29 vials of Novolin were sold.
50-29=21 vials of Humulin were sold.
Proof:
34.39*29+27.06*21=1,565.57
997.31+568.26=1,565.57
1,565.57=1,565.57